# Few-Shot Learner Generalizes Across AI-Generated Image Detection

### v1

@@ Acknowledgments
This work is supported by the National Science and Technology Major Project (No. 2022ZD0116314) and the National Science Foundation of China (No. 62106249).
@@
AI-generated image detection has been a subject of extensive researches and is regarded as an important technology to reduce cheating, misinformation and harm on internet platforms and social media. Our work aims at enhancing fake image detection performance and achieving better generalization with a few samples, rather than increasing synthetic data. All the images used in our experiment are from open-source dataset. We believe our work can contribute to the development of deepfake detecting applications in many real-world scenarios.

### v2

@@ Acknowledgments
This work is supported by the National Science and Technology Major Project (2022ZD0116314), the National Natural Science Foundation of China (62106249, 62476070, 6243000159, 62102416), Shenzhen Science and Technology Program (JCYJ20241202123503005, GXWD20231128103232001), Department of Science and Technology of Guangdong (2024A1515011540), and the Natural Science Foundation of Jiangsu Province under Grant BK20243051.
@@
AI-generated image detection has been a subject of extensive research. It is regarded as an important technology to reduce cheating, misinformation and harm on Internet platforms and social media. This work focuses on enhancing the performance of fake image detection rather than increasing synthetic data. All the images used in our experiments are from open-source datasets. We believe that our work can significantly contribute to the development of deepfake-detection applications in many real-world scenarios.